# Revisiting DDIM Inversion for Controlling Defect Generation by Disentangling the Background

### v1

@@ 4.2 Synthetic Data Quality
For comparing generation quality between ours and other baselines, we measure FID score [8] and LPIPS [27] for each category in MVTec-AD and MVTec-Loco, and use 100 synthetic images per each categories from each generative model trained on split folds in each benchmark datasets. Table 1 demonstrates our methodology can generate more realistic and diverse structure anomaly images, where our method overall outperforms the baseline. In addition, Table 3 shows that our methodology outperforms in logical anomalies, where our model generates the defect considering the background.
@@
Although Anomalydiffusion [10] is also diffusion-based method, it sometimes generates some unrealistic defects depending on the target mask mentioned previously. Figure 3 shows that Anomalydiffusion generates the defects given target mask, but its generation can be located out of objects. Since they usually generate fully filled defects in the masked area, their generation could be low-quality, where its generation can be bias for anomaly detection training. Further generations are shown in Appendix.
@@ 4.3.1 Quantitative Analysis
To demonstrate our synthetic methodology’s effectiveness in anomaly detection, we utilized the generated dataset for training a segmentation backbone. In this experiment, we trained naive U-Net [22] with both real and synthetic defect datasets. We use normal samples in training and 100 defect instances per defect type for 200 epochs followed by the framework of Anomalydiffusion [10].
@@ 4.3.2 Qualitative Analysis
We extend our anomaly detection experiment to analyze the effect of the synthetic datasets from the perspective of the training loss landscape [15], where we visualize the loss landscape of each normal and anomaly training dataset without synthetic instances in Figure. 4. According to several studies [11, 13] indicating that attaining a loss flat minima is closely related to model generalization, our methodology enables more robust learning for unseen anomalies. Especially, loss landscape of anomaly samples is flatter than other baselines, which demonstrates our better performance given anomalies in both image and pixel levels.

### v2

@@ 4.2 Synthetic Data Quality
For comparing generation quality between ours and other baselines, we measure FID score (Heusel et al., 2017) and LPIPS (Zhang et al., 2018) for each category in MVTec-AD and MVTec-Loco, and use 100 synthetic images per each categories from each generative model trained on split folds in each benchmark datasets. Although previous benchmarks do not consider Table 1 demonstrates our methodology can generate more realistic and diverse structure anomaly images, where our method overall outperforms the baseline. In addition, Table 3 shows that our methodology outperforms in logical anomalies, where our model generates the defect considering the background.
@@
Although Anomalydiffusion (Hu et al., 2024) is also diffusion-based method, it sometimes generates some unrealistic defects depending on the target mask mentioned previously. Figure 3 shows that Anomalydiffusion generates the defects given target mask, but its generation can be located out of objects. Since they usually generate fully filled defects in the masked area, their generation could be low-quality, where its generation can be bias for anomaly detection training. Further generations are shown in Appendix A.7.
@@ 4.3.1 Quantitative Analysis
To demonstrate our synthetic methodology’s effectiveness in anomaly detection, we utilized the generated dataset for training a segmentation backbone. In this experiment, we trained naive U-Net (Ronneberger et al., 2015) with both real and synthetic defect datasets. We use normal samples in training and 100 defect instances per defect type for 200 epochs followed by the framework of Anomalydiffusion (Hu et al., 2024).
@@ 4.3.2 Qualitative Analysis
To illustrate how unrealistic synthetic data can compromise the robustness of anomaly detection, we extend our experiments by analyzing the impact of synthetic datasets through the lens of the training loss landscape (Li et al., 2018). Specifically, we visualize the loss landscapes of normal and anomaly training datasets without synthetic instances in Figure. 4. Prior studies (Jiang et al., 2019; Kim et al., 2023) suggest that achieving a flat minimum in the loss function is strongly associated with better model generalization. Following this insight, our approach facilitates more robust learning for detecting unseen anomalies. Since a lower first eigenvalue of the loss Hessian corresponds to a flatter loss landscape, our detection model exhibits superior generalization compared to other baselines, validating its effectiveness in capturing anomalies at both the image and pixel levels.
@@ 5 Conclusion
Appendix A Appendix
A.1 Details of Training
Details of datasets In our experiments on data generation and anomaly detection, we use the MVTec-Ad and MVTec-Loco datasets. Although the training sets for both datasets consist exclusively of normal samples, we establish a supervised setting by splitting the anomaly test set into two folds. Specifically, one fold contains the first half of the anomaly test samples, while the other fold contains the second half. During training, we first train the diffusion model using one fold of anomalies and perform inference using the normal samples from the training set. Next, we train the U-Net using the original normal training samples, the anomalies from one fold, and synthetic anomalies. For evaluation, we assess the model’s performance on the remaining fold of anomalies and the normal test samples. For the unsupervised-based methods, we just utilize normal samples for training, which are identical settings with their paper.
Details of training diffusion model In our framework, we utilize the stable diffusion with $256\times 256$ resolutions. We train the only U-Net part with 500 iterations per each class of objects.
Details of Inference After training the diffusion model with one-fold anomalies, we utilize the DDIM inversion with 50 time steps to train normal samples and generate anomalies with a ground-truth mask in one-fold.
A.2 Proof of Lemma 3.1
Lemma A.1.
$(1-m)\odot z_{0}=(1-m)\odot\sqrt{1\over\alpha_{t}}\left[z_{t}^{m}-\sqrt{1-\alpha_{t}}\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2})\right]$
Proof.
$\mathcal{L}(\theta^{*})$ $=||\epsilon-m\odot\epsilon_{\theta^{*}}(z_{t},t,C_{1})-(1-m)\odot\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2}))||_{2}^{2}$
$=0$ (14)
$\epsilon$ $=m\odot\epsilon_{\theta^{*}}(z_{t},t,C_{1})+(1-m)\odot\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2}))$ (15)
From Eq.6, we suppose that the optimal denoised equation of $z_{t}$ as follows.
$z_{0}$ $=\sqrt{1\over\alpha_{t}}\left[z_{t}-\sqrt{1-\alpha_{t}}\epsilon\right]$ (16)
$=\sqrt{1\over\alpha_{t}}\left[z_{t}-\sqrt{1-\alpha_{t}}(m\odot\epsilon_{\theta^{*}}(z_{t},t,C_{1})+(1-m)\odot\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2}))\right]$ (17)
Due to $z_{t}^{m}=(1-m)\odot z_{t}\ \forall t$, we prove this statement.
$(1-m)\odot z_{0}$ $=(1-m)\odot\sqrt{1\over\alpha_{t}}\left[z_{t}-\sqrt{1-\alpha_{t}}\epsilon\right]$ (18)
$=\sqrt{1\over\alpha_{t}}\left[(1-m)\odot z_{t}-\sqrt{1-\alpha_{t}}((1-m)\odot\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2}))\right]$ (19)
$=\sqrt{1\over\alpha_{t}}\left[z_{t}^{m}-\sqrt{1-\alpha_{t}}((1-m)\odot\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2}))\right]$ (20)
A.3 Proof of Theorem 3.2
Theorem A.2.
Suppose that $\mathcal{L}(\theta^{*})=0$ and $\alpha_{t}<\alpha_{t-1}$ for all $t$ , then $\lim_{\alpha_{t}\rightarrow 0}\tilde{z_{t}}-\tilde{z_{t}}^{m}=m\odot\epsilon$, $\epsilon\sim\mathcal{N}(0,I)$
Proof.
Suppose that $\mathcal{L}(\theta^{*})=0$ and we reformulate DDIM sampling as Euler method followed by (Song et al., 2021a).
${z_{t-\Delta t}\over\sqrt{\alpha_{t-\Delta t}}}={z_{t}\over\sqrt{\alpha_{t}}}+\left(\sqrt{{1-\alpha_{t-\Delta t}\over\alpha_{t-\Delta t}}}-\sqrt{{1-\alpha_{t}\over\alpha_{t}}}\right)\epsilon_{\theta^{*}}(z_{t},t,C)$ (21)
For $\Delta t\rightarrow 0$, we inverse the DDIM Sampling as follows:
$z_{t}=\sqrt{\alpha_{t}}[\sqrt{1\over\alpha_{t-1}}z_{t-1}+(\sqrt{{1\over\alpha_{t}}-1}-\sqrt{{1\over\alpha_{t-1}}-1})\cdot\epsilon_{\theta^{*}}(z_{t-1},t-1,C)$ (22)
$\tilde{z}_{t}$ is forwarded by DDIM Inversion as follows:
$\tilde{z}_{t}=\sqrt{\alpha_{t}}[\sqrt{1\over\alpha_{t-1}}\tilde{z}_{t-1}+(\sqrt{{1\over\alpha_{t}}-1}-\sqrt{{1\over\alpha_{t-1}}-1})\cdot((1-m)\odot\epsilon_{\theta^{*}}(\tilde{z}_{t-1}^{m},t-1,C_{2})+m\odot\epsilon_{\theta^{*}}(\tilde{z}_{t-1},t-1,C_{1}))]$ (23)
From Theorem.1 of (Song et al., 2021a), $m\odot\epsilon_{\theta^{*}}(\tilde{z}_{t-1},t-1,C_{1})=m\odot\epsilon$ is satisfied, then the following equations are derived.
$\tilde{z}_{t}$ $=\sqrt{\alpha_{t}}[\sqrt{1\over\alpha_{t-1}}\tilde{z}_{t-1}+(\sqrt{{1\over\alpha_{t}}-1}-\sqrt{{1\over\alpha_{t-1}}-1})\cdot((1-m)\odot\epsilon_{\theta}(\tilde{z}_{t-1}^{m},t-1,C_{2})+m\odot\epsilon)]$ (24)
$\tilde{z}_{t}-\tilde{z}_{t}^{m}$ $=\sqrt{\alpha_{t}}[\sqrt{1\over\alpha_{t-1}}m\odot\tilde{z}_{t-1}+(\sqrt{{1\over\alpha_{t}}-1}-\sqrt{{1\over\alpha_{t-1}}-1})\cdot(m\odot\epsilon)]$ (25)
Therefore, $\lim_{\alpha_{t}\rightarrow 0}\tilde{z_{t}}-\tilde{z_{t}}^{m}=m\odot\epsilon$
A.4 Proof of Proposition 3.3
Proposition A.3.
Proof.
Due to $z_{T}^{*}=z_{T}^{m}+m\odot\epsilon$ and $(1-m)\odot z_{T}^{*}=z_{T}^{m}$
$z_{T-1}^{m}$ $=\sqrt{\alpha_{T-1}\over\alpha_{T}}z_{T}^{m}+(1-m)\odot(\sqrt{{1\over\alpha_{T-1}}-1}-\sqrt{{1\over\alpha_{T}}-1})\cdot\epsilon_{\theta}(z_{T}^{m},T,C_{2})$ (26)
$=(1-m)\odot\sqrt{\alpha_{T-1}\over\alpha_{T}}z_{T}^{*}+(1-m)\odot(\sqrt{{1\over\alpha_{T-1}}-1}-\sqrt{{1\over\alpha_{T}}-1})\cdot\epsilon_{\theta}(z_{T}^{m},T,C_{2})$ (27)
From Eq.3.2, we denoise $z_{T}^{*}$ as follows.
$z_{T-1}^{*}=\sqrt{\alpha_{T-1}\over\alpha_{T}}z_{T}^{*}+(\sqrt{{1\over\alpha_{T-1}}-1}-\sqrt{{1\over\alpha_{T}}-1})\cdot((1-m)\odot\epsilon_{\theta}(z_{T}^{m},T,C_{2})+m\odot\epsilon_{\theta}(z_{T}^{*},T,C_{1}))$ (28)
$(1-m)\odot z_{T-1}^{*}=\sqrt{\alpha_{T-1}\over\alpha_{T}}(1-m)\odot z_{T}^{*}+(\sqrt{{1\over\alpha_{T-1}}-1}-\sqrt{{1\over\alpha_{T}}-1})\cdot((1-m)\odot\epsilon_{\theta}(z_{T}^{m},T,C_{2}))$ (29)
Then, $(1-m)\odot z_{T-1}^{*}=z_{T-1}^{m}$. For arbitrary $k\leq T$, we also prove following equality.
$(1-m)\odot z_{T-k-1}^{*}$ $=\sqrt{\alpha_{T-k-1}\over\alpha_{T-k}}z_{T-k}^{*}+(\sqrt{{1\over\alpha_{T-k-1}}-1}-\sqrt{{1\over\alpha_{T-k}}-1})\cdot((1-m)\odot\epsilon_{\theta}(z_{T-k}^{m},T-k,C_{2}))$ (30)
Suppose $t=T-k-1$, then $||(1-m)\odot(z_{t}^{m}-\tilde{z}_{t}^{m})||_{2}^{2}=||(1-m)\odot(z_{t}^{*}-\tilde{z}_{t}^{m})||_{2}^{2}$
A.5 Details of attention mechanism in Ours
As mentioned before, U-Net is constructed with self-attention and cross-attention layers. In this section, we elaborate on each mechanism with linear layers.
Firstly, we mask the cross-attention with a resized mask, $m^{\prime}$. Spatial feature from the latent $z_{t}$ is transformed to the Query vector, $Q=\ell_{Q}(\phi(z_{t}))$, where $\ell_{Q}$ is linear projection and $\phi$ is spatial feature network of latent. Similarly, we extend key and value vector; $[K_{def},K_{bg}]=[\ell_{K}(C_{def}),\ell_{K}(C_{bg})]$ and $[V_{def},V_{bg}]=[\ell_{V}(C_{def}),\ell_{V}(C_{bg})]$. From the definition, we define the output of cross-attention with $z_{t}$ as follows:
$Attn^{C}(z_{t},C_{1})=m^{\prime}\odot Softmax({QK_{def}^{T}\over\tau})V_{def}+(1-m^{\prime})\odot Softmax({QK_{bg}^{T}\over\tau})V_{bg}$ (31)
Secondly, we define the self-attention, where we do not directly mask the self-attention. Spatial feature from the masked latent $z_{t}^{m}$ is transformed to the Query vector, $\hat{Q}^{m}=\ell_{\hat{Q}}(\hat{\phi}(z_{t}^{m}))$, where $\ell_{\hat{Q}}$ is linear projection and $\hat{\phi}$ is spatial feature network of latent in self-attention layer. Similarly, we extend key and value vector like cross-attention layers, and we define the output of self-attention as follows:
$Attn^{S}(z_{t}^{m})$ $=Softmax({\hat{Q}^{m}\hat{K}^{mT}\over\tau})\hat{V}^{m}$
$=Softmax({\hat{Q}^{m}\hat{K}^{mT}\over\tau})(1-m^{\prime})\odot\hat{V}^{m}$
$=(1-m^{\prime})\odot Attn^{S}(z_{t}^{m})$ (32)
A.6 Details of Qualitative Analysis
To qualitatively demonstrate performance gain of anomaly detection given our synthetic dataset, we conduct visualizing the loss landscapes given train real anomaly and normal instances. Given trained detection model parameter $\theta_{det}$, the loss landscapes as follows:
$g(\alpha,\beta)={1\over n}\sum_{i=1}^{n}l(x_{i},y_{i};\theta_{det}+\alpha d_{1}+\beta d_{2})$ (33)
Followed by Anomaly diffusion (Hu et al., 2024), we utilize the Focal Loss, $l$ with random direction $d_{1}$,$d_{2}$ and scaling scalar value $\alpha$, $\beta$.
A.7 Additional Defect Synthetic Data
Figure 5: Additional defect generation of DFMGAN in MVTec-Ad
Figure 6: Additional defect generation of Anomalydiffusion in MVTec-Ad
Figure 7: Additional defect generation of Ours in MVTec-Ad
Figure 8: Additional defect generation of DFMGAN in MVTec-Loco
Figure 9: Additional defect generation of Anomalydiffusion in MVTec-Loco
Figure 10: Additional defect generation of Ours in MVTec-Loco